# Bridging Today and the Future of Humanity: AI Safety in 2024 and Beyond

### v3

@@ 4 Bridging Today and the Future
This section outlines an ambitious roadmap to the advanced future, highlighting key milestones in the evolution of AI and human civilization. By exploring potential paradigm shifts in fundamental theories and technological breakthroughs, this section categorizes this journey into three eras, i.e., i) the data-driven era, in which AI learns from vast training datasets, ii) the neural augmentation era, where human-machine interface [151, 84, 262] and intelligent chips [89, 119] expand the boundaries of human cognition, and iii) the revolutionary era, characterized by a breakthrough in energy production that catalyzes transformative changes across all aspects of human society; see §4.1, §4.2, and §4.3, respectively. We then examine the evolving paradigms of AI safety throughout this journey, analyzing the focuses and challenges of each era while exploring the impacts of technological advancement on AI safety in §4.4.
@@ 4.1 Rookie Level: Data-Driven Era
The data-driven era represents the earliest phase of AI development, where AI models rely heavily on the quality, availability, and quantity of training data [135, 261, 269]. AI makes probabilistic inferences based on patterns found in massive training datasets, thus failing to achieve genuine understanding of data [100, 245]. While AI in this stage cannot be regarded as real intelligence, it can still assist people to solve problems and enhances productivity in human society.
Stage 1: Past, Current, and the Near Future. This stage is characterized by developing diverse AI models to process various forms of data across different scenarios [135, 53, 66, 215, 236, 268]. Researchers have been engaging in refining model architectures and developing advanced algorithms, enabling more sophisticated analysis of data in different formats. The thrive of LLMs exemplifies this progress, demonstrating unprecedented capabilities of AI models in language processing and generation [268, 297, 236]. Concurrently, significant advancements have also been observed in other domains such as computer vision [105, 226, 264], image processing [241, 25, 221], video analysis [228, 270, 276], and audio recognition [67, 182, 4, 110]. This stage is driven by the exploration of vast datasets and the continuous refinement of algorithms, which consistently pushes the boundaries of AI capabilities across specific domains.
Stage 2: Interdisciplinary AI. This stage is characterized by the integration of AI through cross-disciplinary applications, driving significant innovations and transformations in traditional fields such as healthcare [236, 2], finance [268, 292], and education [188, 139]. In this stage, AI applications would unlock new opportunities across industries and reshape human society. Mature multi-agent systems [120, 253, 272, 82] that leverage AI entities with specialized capabilities might be widely deployed to assist people to solve sophisticated problems. Such agents will participate in decision-making processes or represent users to complete some tasks. For example, in healthcare, a multi-agent system may involve collaborations between multiple specialized medical agents and integrate data from various sources to formulate comprehensive treatment plans. The rise of interdisciplinary AI integration will enhance AI’s utility in increasingly complex scenarios. While AI at this stage does not yet achieve true intelligence, it becomes deeply integrated across various domains in human society and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
@@ 4.2 Champion Level: Neural Augmentation Era
In the neural augmentation era, humans may witness breakthroughs in fundamental theories and advanced technologies, which may extend the boundary of human cognition and enhance life experience.
Stage 3: Breakthroughs in Advanced Technologies. At this stage, breakthroughs in advanced technologies, such as brain-computer interfaces [84, 151, 262], nanotechnology [140, 19, 64], and holographic technology [22, 283, 15], augment human senses and cognition, enabling unprecedented levels of interactions between humans and machines. Language barriers may disappear, as human can perceive with chips for understanding, which enables ubiquitous real-time context-aware translations between people of different languages and cultures. Intelligent chips [89, 119] might be integrated into human bodies, creating a seamless interface between biological and digital systems. Such advancements may change the nature of the data. Unlike data of static formats such as text, images, or videos, the data might be more comprehensive, in dynamic and sensory-rich formats, such as neural activity patterns or real-time environmental data from augmented reality systems [31, 32]. This stage expands the boundaries of human perception and cognition, revolutionizing how people experience and interact with the world. However, applying these techniques will be still challenging due to the energy constraints and the significant resources required by these technologies.
Stage 4: Breakthroughs in Fundamental Theories. At this stage, humanity may witness a new wave of breakthroughs in fundamental theories, such as Einstein’s theories of relativity [60, 61] and quantum mechanics [62, 289]. These theoretical advancements have the potential to redefine our understanding of the universe, and, moreover, offer new possibilities for groundbreaking technological innovations and provide critical insights into solving the longstanding challenges in human history, i.e., the energy issues.
@@ 4.3 Ultimate Level: Revolutionary Era
Stage 5: Energy Revolution. A new era would arrive as humanity finally achieves energy revolution, unraveling the mysteries of sustainable power generation. This may be achieved through mature technologies that are currently on the horizon, such as breakthroughs in large-scale controlled nuclear fusion [73, 205, 187], or revolutionary solar technologies [23, 141], together with advancement in nanotechnologies [140, 19, 64] that produce new materials for collecting and storing solar energy efficiently. The energy issue may also be solved through innovations that are beyond the scope of today’s technologies, such as deploying satellites for gathering and storing solar energy in space. These satellites might then transmit the abundant energy to the receivers on Earth. Regardless of the specific technology, the flood of clean, abundant energy can eliminate our dependence on fossil fuels, and revolutionize the global economic landscapes as well as human civilization.
Stage 6: The Internet of Everything. The final stage is the Internet of Everything. Every object, living being, and system become parts of a vast, intelligent network. This hyper-connected world is characterized by unprecedented levels of perception, communication, and interactions between humans and intelligent robots. In this era, everyone is connected to a huge network, and the boundary of human sense has been extended, which may also catalyze breakthroughs in holographic technology [15, 22, 283]. Remote experiences become true; working, entertainment, and communication are revolutionized, and human has entered the ultimate advanced era.
@@ 4.4 Evolving AI Safety Paradigms
The landscape of AI safety shifts as humanity processes through stages of technology revolutions, and the focus of AI safety is different at each stage. The core AI safety challenges in different historical eras are summarized in Figure 3.
@@
Data-Driven Era. Current AI safety practices primarily focus on data safety and model safety [39, 106, 44, 225, 171, 198, 260, 230, 229]. AI safety researchers and practitioners have explored adversarial attacks targeting training data and AI models [9, 36, 123, 250, 216, 279] to achieve malicious goals, for example, designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection attacks [78, 303, 130, 34, 93, 266, 304]. Correspondingly, red-teaming ML models [12, 71, 177, 202, 249] is widely used to identify potential vulnerabilities in AI models, and significant work has been proposed to enhance security and robustness of AI systems [27, 246, 301], aligning AI with human values [256, 178, 237, 211, 252]. Data privacy is also a critical aspect, given the vast amounts of sensitive information in the training data [122, 124, 86, 121]. Additionally, researchers have been engaging in AI governance to enhance AI accountability [21], transparency and explainability [302] in AI systems to ensure the behaviors of AI are understandable, accountable, and aligned with regulations [21, 136, 219]. .
@@
Neural Augmentation Era. Cross-disciplinary AI applications might have been widely deployed in human society, and advancements in intelligent chips [89, 119] and brain-computer interfaces [84, 151, 262] might extend the boundaries of human cognition and have changed the way of human perceiving the world. With the integration of intelligent chips and their neural connections to intelligent systems, users can interact with machines and AI with high flexibility, and the customization of AI services for different user needs might be a critical focus [255]. Individuals might rely on AI in different ways, and personalized experiences of AI services might be explored based on users’ background and requirements.
@@
Revolutionary Era. In the revolutionary era, AI safety challenges may focus on complex interactions between interconnected entities, e.g., humans and intelligent robots, within the network. Robots might provide customized services according to different user requirements, and ensuring the safety and reliability of the interactions between them becomes challenging. Robots may need to integrate data from external sources in real-time to fulfill user requests. Managing the quality and security of these interactions is crucial, as any breakdown or interference could lead to incorrect or even harmful outcomes. Thus, ensuring seamless communication, data integrity, and efficient decision-making in such a sophisticated network may require smarter and more efficient methods for managing, such as advanced cache management systems, to support the dynamic and complex interactions between humans and robots in the network.
@@
By aligning advancements of AI with human civilization, we can see the focuses of AI safety evolve in different stages. Initially, the focus is on addressing safety issues related to data and models. With progress of AI and advanced technologies, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs. Finally, when humanity entering the era of Internet of Everything, challenges in AI safety further evolve, focusing on complex interactions between humans and robots in the network. These insights offer guidelines for our efforts on AI safety at the current stage. The following sections will overview today’s AI safety landscape, then reflect on the progress in this area so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed at the current stage.
5 Back to the 2020s: Today’s AI Safety Landscape
This section overviews current mainstream AI safety practices, and project them against the blueprint future in §3.1 to guide our current efforts on AI safety.
5.1 Jailbreaking and Red Teaming
Jailbreaking [51, 127, 179, 212, 220, 224, 240, 257, 266, 277, 304] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs. Malicious users or adversaries can exploit vulnerabilities in AI models through crafted inputs or backdoors, inducing them to produce inappropriate responses, such as giving instructions for illegal activities or generating offensive content. These activities pose significant risks by compromising safety measures of AI models, making it challenging to ensure the model to behave appropriately, especially in real-world scenarios where the potential inputs are unpredictable.
As AI systems grow more sophisticated, red-teaming must continually adapt to counter increasingly complex attacks, which calls for more advanced tools and interdisciplinary expertise [157]. Model owners can conduct diverse red teaming regularly to anticipate and mitigate risks in evolving use cases. Also, current red-teaming practices are often static, typically conducted after model pretraining with one-time attacks. Thus, these practices potentially fail to address evolving user needs and complex interactions between users and models, especially when dealing with customized AI services for different user groups. To address this, red teaming should involve more diverse use cases, such as employing different user groups and simulating complex interactions with dynamic contexts, to address safety risks in customized AI services and enable a comprehensive evaluation of potential vulnerabilities across different scenarios.
5.2 Content Moderation and Customized Moderation Pipeline
Content moderation involves detecting and reviewing both user input and AI-generated content to ensure compliance with safety, ethical, and regulatory standards [137, 116, 83, 94]. Moderation approaches employ rule-based methods, machine learning classifiers, and human oversight to monitor, evaluate, and manage the inputs and outputs of AI models. These approaches identify inappropriate content, such as toxicity, bias, hallucinations, private information, and jailbreaks [251]. By preventing AI models from engaging with or producing harmful content, content moderation plays a critical role in maintaining responsible AI behavior. The customized moderation pipeline [145, 194] employs multiple components at different stages of LLM inference to enhance the quality of the model outputs while enhancing the control of the whole AI workflows. It functions as a flexible intermediate layer between users and AI models, enabling users to add customized components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
As we reflect on practices for AI moderation and customized moderation pipelines, several key challenges and considerations have emerged. 1) Hallucination remains a significant issue and is unlikely to solve completely, regardless of how AI evolves. That is because hallucinations stem from nearly every stage of the LLM lifecycle, including training data quality, model architecture, and inference processes, etc [92]. Despite advancements in AI technology, hallucinations persist as an inherent challenge that requires constant attention in moderation practices. 2) The balance between utility and safety is crucial and may require different solutions according to the changing contexts and user needs. Overly strict moderation can cause an AI system to be helpless, e.g., an LLM that always responds “sorry, I cannot answer this question” is completely safe but useless. Defining appropriate thresholds for acceptable responses is challenging and requires considering potential risks and the context of user queries. 3) Customization of moderation methods should be applied for different use cases, as certain information may be safe for one user group but inappropriate for another. As an example, a chemistry industry practitioners may need specific and technical information on dangerous substances like explosives for professional purposes, while such information should be restricted for the general public. Current moderation practices might fail to include such cases, and applying identical restrictions between different user expertise or context might limit the utility of AI in specialized fields. 4) Some moderation measures, such as detection of toxicity, bias, stereotypes, out-of-distribution content, might be unnecessary for most real-world use cases. This is because most of the current AI applications and multi-agent systems [120, 37] leverage existing LLM APIs [139, 188, 137], open-source LLMs [239, 278, 97], or utilizing opensource LLMs as base models, many of which already have built-in components or have done safety alignments for managing inappropriate inputs and outputs. Incorporating certain detection mechanisms may offer limited benefits to the safety of LLM inputs and outputs while increasing unnecessary computational overhead.
@@ 5.2 Content Moderation and Customized Moderation Pipeline
5.3 AI Privacy

### v4

@@ 4 Bridging Today and the Future
This section outlines an ambitious roadmap to the advanced future, highlighting key milestones in the evolution of AI and human civilization. By exploring potential paradigm shifts in fundamental theories and technological breakthroughs, this section categorizes this journey into three eras, i.e., i) the data-driven era, in which AI learns from vast training datasets (§4.1), ii) the neural augmentation era, where human-machine interface [194, 97, 319] and intelligent chips [104, 149] expand the boundaries of human cognition (§4.2), and iii) the revolutionary era, characterized by a breakthrough in energy generation that transforms human society (§4.3). We then examine the evolving paradigms of AI safety and analyze our focuses and challenges throughout this journey (§4.4).
@@ 4.1 Rookie Level: Data-Driven Era
The data-driven era represents the earliest phase of AI development, where AI models rely heavily on the quality, availability, and quantity of training data [173, 318, 359, 327]. In this era, AI makes probabilistic inferences based on patterns found in the training datasets, thus failing to achieve genuine understanding of data [123, 300]. While AI in this stage cannot be regarded as real intelligence, it can still assist people to solve problems and enhances productivity in human society.
Stage 1: Past, Current, and the Near Future. This stage is characterized by developing diverse AI models for different tasks [173, 60, 75, 92, 325, 231, 268, 290, 362, 326]. Researchers have been engaging in refining model architectures and developing advanced algorithms and computational techniques, enabling more sophisticated analysis of data in different formats for various tasks [91, 156, 348, 60, 75, 325, 231, 172, 92, 85, 49]. The thrive of LLMs exemplifies this progress, demonstrating unprecedented capabilities of AI models in language processing and generation [326, 360, 290]. Significant advancements have also been observed in other domains such as computer vision [129, 279, 321, 311, 61], image processing [296, 28, 274], video analysis [281, 328, 334], audio recognition [76, 226, 4, 135], etc.
Stage 2: Interdisciplinary AI. This stage is characterized by the integration of AI through cross-disciplinary applications, leading to innovations in traditional fields such as healthcare [290, 2], finance [326, 354], and education [235, 178]. AI applications would unlock new opportunities across industries and reshape human society. Mature multi-agent systems [150, 309, 330, 94] that leverage AI entities with specialized capabilities might be widely deployed to assist people to solve sophisticated problems. Such agents will participate in decision-making processes or represent users to complete some tasks. For example, in healthcare, a multi-agent system may involve collaborations between multiple specialized medical agents and integrate data from various sources to formulate comprehensive treatment plans. The rise of interdisciplinary AI integration will enhance AI’s utility in increasingly complex scenarios. While AI at this stage does not yet achieve real intelligence, it becomes deeply integrated across various domains in human society and may catalyze the emergence of new fields at the intersection of traditional disciplines and AI.
@@ 4.2 Champion Level: Neural Augmentation Era
In this era, humans may witness breakthroughs in fundamental theories and advanced technologies, which may extend the boundary of human cognition and enhance life experience.
Stage 3: Breakthroughs in Advanced Technologies. Breakthroughs in advanced technologies, such as brain-computer interfaces [97, 194, 319], nanotechnology [179, 22, 73], and holographic technology [25, 342, 18], augment human cognition and enable unprecedented levels of interactions between humans and machines. Language barriers may disappear, as human can perceive with chips for understanding, enabling ubiquitous real-time context-aware translations between people of different languages and backgrounds. Intelligent chips [104, 149] might be integrated into human bodies, creating a seamless interface between biological and digital systems. Such advancements may fundamentally transform the nature of data. Unlike traditional static formats such as text, images, or videos, future data is likely to become more comprehensive, dynamic, and sensory-rich, such as neural activity patterns or real-time environmental data collected from augmented reality systems [34, 35]. This stage expands the boundaries of human perception and cognition, revolutionizing how people experiencing and interacting with the world. However, applying these techniques might be still challenging due to the significant resources required by these technologies.
Stage 4: Breakthroughs in Fundamental Theories. At this stage, humanity may witness breakthroughs in fundamental theories, such as Einstein’s theories of relativity [69, 70] and quantum mechanics [71, 351]. Such advancements have the potential to redefine our understanding of the universe, and, moreover, offer new possibilities for groundbreaking technological innovations and provide critical insights into solving the longstanding challenges in human history, i.e., the energy issue.
@@ 4.3 Ultimate Level: Revolutionary Era
Breakthroughs in fundamental theories and advanced technologies provide building blocks for addressing the energy issue, unlocking new possibilities for productivity and innovation, and ultimately achieving the Internet of Everything.
Stage 5: Energy Revolution. Humanity secures a new paradigm with the energy revolution that eliminates reliance on fossil fuels, which ensures environmental sustainability and redefines global economic ecosystems. This milestone may be achieved through several potential technical pathways.
i) Controlled Nuclear Fusion. Significant challenges of nuclear fusion include managing high-pressure plasma environments and developing materials that can withstand extreme temperatures and intense radiation [13, 186, 110, 14]. Theoretical and technological breakthroughs in physics [183, 90, 253] and materials science [179, 22, 73] may overcome these obstacles and enable widespread deployment of small ion thrusters [126, 59, 225], which transforms propulsion and energy transmission by providing abundant clean energy.
ii) Revolutionary Solar Technologies. Advancements in solar technologies [26, 180] and nanotechnology [179, 22, 73] might unlock unprecedented methods for efficiently collecting and storing solar energy, e.g., with lightweight solar panels and flexible storage solutions.
iii) Emerging New Technologies. Novel technologies for energy generation beyond today’s imagination might be developed in the future. One possibility is to deploy of satellites for collecting and storing solar energy in space [133, 41, 100, 302], where the vacuum of space eliminates energy reflection or absorption by the atmosphere, allowing satellites to convert and transmit solar energy to receivers on Earth efficiently.
Regardless of the specific technological path, the flood of clean and abundant energy would eliminate our dependence on fossil fuels, reshaping the global economies and elevating human civilization to unprecedented levels of prosperity and innovation.
Stage 6: The Internet of Everything. The ultimate stage is the Internet of Everything, where every object, every living being, and every system integrate into a vast network. This hyper-connected net involves unprecedented levels of complex interactions, perceptions, and communications between humans and intelligent objects, e.g., robots. Moreover, holographic technology [18, 25, 342] will significantly enhance human experience in the network by making remote interactions as vivid as physical presence. In this stage, interactions between humans and the world are re-defined, marking a new chapter in human civilization.
@@ 4.4 Evolving AI Safety Paradigms
The AI safety landscape evolves as humanity progresses through various stages of scientific and technological advancements, with the focus of AI safety differing at each stage. The primary AI safety challenges across different historical eras are summarized in Figure 3.
@@
Data-Driven Era. AI safety practices primarily focus on data safety and model safety [44, 130, 50, 278, 214, 247, 317, 283, 282]. AI safety researchers and practitioners have explored adversarial attacks targeting training data and models [10, 39, 154, 306, 269, 337] to achieve malicious goals, e.g., designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection [88, 367, 166, 37, 108, 323, Zou2023UniversalAT]. Correspondingly, red-teaming is widely used to identify potential vulnerabilities in AI models [15, 80, 220, 252, 305], while significant work has been proposed to address safety risks and enhance security and robustness of AI systems [138, 324, 232, 72, 30, 301, 365], align AI with human values [313, 221, 292, 264, 308], etc. Data privacy is also a critical aspect, given the vast amounts of sensitive information in the training data [153, 155, 101, 151]. Moreover, researchers have been engaging in AI accountability [24], transparency and explainability [366] in AI systems to ensure the behaviors of AI are understandable, accountable, and aligned with regulations [24, 175, 272].
@@
Neural Augmentation Era. Cross-disciplinary AI applications might have been widely deployed in human society, and advancements in intelligent chips [104, 149] and brain-computer interfaces [97, 194, 319] might extend the boundaries of human cognition. With the integration of intelligent chips and their neural connections to intelligent systems, users with different needs can interact with machines and AI with high flexibility, and the customization of AI services for different users might be a focus [312]. Individuals might rely on AI in different ways, and personalized experiences of AI services might be explored based on users’ background and requirements.
@@
Revolutionary Era. When the Internet of Everything becomes a reality, AI safety challenges may focus on complex interactions between interconnected entities, e.g., humans and intelligent robots, within the network. Robots might provide customized services according to different user requirements, thus, ensuring safety and reliability of the interactions between them becomes challenging. Robots may need to integrate data from external sources in real-time to fulfill user requests. Managing the quality and security of these interactions is crucial, as any breakdown or interference could lead to incorrect or even harmful outcomes. Thus, ensuring high-quality communication, data integrity, and efficient decision-making in such a sophisticated network may require smart and efficient management methods, such as advanced cache management systems, to support the dynamic and complex interactions between humans and robots in the network.
@@
By aligning advancements of AI with human civilization, we can see the focuses of AI safety evolve in different stages. Initially, the focus is on addressing safety issues related to data and models. With advancements of AI and advanced technologies, the focus of AI safety shifts to customization, which ensures the services provided by AI systems to meet individual needs. Finally, when humanity entering the era of Internet of Everything, challenges in AI safety further evolve, focusing on complex interactions between humans and robots in the network. These insights offer guidelines for our efforts on AI safety at the current stage. The following sections will overview today’s AI safety landscape, then reflect on the progress in this field so far, and finally, explore the challenges and responsibilities of current AI safety practitioners, highlighting areas where further improvements and efforts are needed at the current stage.
@@
5 Back to the 2020s: Today’s AI Safety Landscape
This section overviews current mainstream AI safety practices, and project them against the blueprint future to guide our current efforts on AI safety. Table 4 summarizes research and industry efforts in each safety topic.
5.1 Jailbreaking and Red Teaming
Jailbreaking [57, 163, 222, 265, 273, 277, 295, 314, 323, 335, Zou2023UniversalAT] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs. Malicious users or adversaries can exploit vulnerabilities in AI models through crafted inputs or backdoors, inducing them to produce inappropriate responses, such as giving instructions for illegal activities or generating offensive content [57, 163, 222, 265, 277, 295, 335]. These activities pose risks by compromising safety measures of AI models, making it challenging to ensure the model to behave appropriately, especially in real-world scenarios where the potential inputs are unpredictable.
As AI systems grow sophisticated, red-teaming must continually adapt to counter increasingly complex attacks, calling for more advanced tools and interdisciplinary expertise [200]. Also, current red-teaming practices are often static, typically conducted after model pretraining with one-time attacks. Thus, these practices potentially fail to address evolving user needs and complex interactions between users and models, especially when dealing with customized AI services for different user groups. To address these problems, model owners can conduct diverse red teaming regularly to anticipate and mitigate risks in evolving use cases. Red teaming should also involve more diverse use cases, such as employing different user groups and simulating complex interactions with dynamic contexts, to address safety risks in customized AI services and enable a comprehensive evaluation of potential vulnerabilities across different scenarios.
5.2 Content Moderation and Customized Moderation Pipeline
Content moderation employs rule-based methods, machine learning classifiers, and human oversight to detect and review user inputs and AI-generated content to ensure compliance with safety, ethical, and regulatory standards [176, 146, 95, 111]. These approaches identify inappropriate content, including toxicity, bias, hallucinations, private information, and jailbreaks, etc [307]. By preventing AI models from engaging with or producing harmful content, content moderation plays a critical role in maintaining responsible AI behavior.
Customized moderation pipelines [185, 243] employ multiple components at different stages of LLM inference to enhance the quality of the model outputs while enhancing the control of the whole AI workflows. They function as a flexible intermediate layer between users and AI models, enabling users to add customized components, such as code-based rules and quality guarantees, at any stage of inference based on their personalized needs.
As we reflect on practices for AI moderation and customized moderation pipelines, several key challenges and considerations have emerged. 1) Hallucination remains a significant issue and is unlikely to solve completely, regardless of how AI evolves. That is because hallucinations stem from nearly every stage of the LLM lifecycle, including training data quality, model architecture, and inference processes, etc [107]. Despite advancements in AI technology, hallucinations persist as an inherent challenge that requires constant attention in moderation practices. 2) The balance between utility and safety is crucial and may require different solutions according to the changing contexts and user needs. Overly strict moderation can cause AI models to be uninformative, e.g., an LLM that always responds “sorry, I cannot answer this question” is completely safe but useless. Defining appropriate thresholds for acceptable responses is challenging and requires considering potential risks and the context of user queries. 3) Customization of moderation methods should be applied for different use cases, as certain information may be safe for one user group but inappropriate for another. As an example, a chemistry industry practitioners may need specific and technical information on dangerous substances like explosives for professional purposes, while such information should be restricted for the general public. Current moderation practices might fail to include such cases, and applying identical restrictions between different user expertise or context might limit the utility of AI in specialized fields. 4) Some moderation measures, such as detection of toxicity, bias, stereotypes, out-of-distribution content, might be redundant for most real-world use cases. This is because most of the current AI applications and multi-agent systems [150, 40] leverage existing LLM APIs [178, 235, 176], open-source LLMs [294, 336, 116], or utilize open-source LLMs as base models for fine-tuning, many of which already have built-in safety components or have done safety alignments for handling inappropriate content. Incorporating certain detection mechanisms may offer limited benefits to the safety of model inputs and outputs while increasing unnecessary computational overhead.
5.3 Privacy